OpenStudy (anonymous):

The figures below show two different ways of arranging four identical triangles of grey poster board on top of a white square. The square has sides equal to x + y, while the hypotenuse of each triangle is represented by the variable c.

OpenStudy (anonymous):

Hazel wrote the following statements to prove that c2 = x2 + y2. 1. Area of the four grey triangles inside figure A = 2. Area of the white square inside figure A = c2 3. Area of figure A = c2 + 2xy 4. Area of the four grey triangles inside figure B = 5. Area of the two white squares inside figure B = x2+ y2 6. Area of figure B = 2x2+ 2y2 + 2xy 7. Area of figure A = area of Figure B, hence c2 + 2xy = 2x2+ 2y2 + 2xy 8. Therefore, c2 = x2+ y2

OpenStudy (anonymous):

Looks good. What's the problem?

OpenStudy (anonymous):

They want to know which is the first incorrect statement in Hazel's Proof

OpenStudy (anonymous):

Statement 4 is wrong

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

your welcome